OpenStudy (anonymous):

Riddle me this , can somebody show me step by step on how to get 10/3 , from a over b+c if a/b=5;b/c=2

OpenStudy (amistre64):

maybe restructure it so that everything speaks in terms of b

OpenStudy (asnaseer):

you've already asked this question here: http://openstudy.com/users/thx1138#/updates/4fb80862e4b05565342d6e40 if the explanations were not clear then you should have asked for more explanation in that question rather than closing it and reposting the same question.

OpenStudy (amistre64):

yes, that would have been the best thing to do.

OpenStudy (anonymous):

Then can one of you please clear it up for me?

OpenStudy (amistre64):

i cant read your mind so i dont know how this can NOT possibly be clear as it stands :/ you are going to have to be more specific in your perplexities

OpenStudy (asnaseer):

this was the explanation given by @rs32623 in the other question: from a/b=5 we have a=5b from b/c=2 we have c=b/2 now for a/(b+c) we have 5b/(b+b/2)=10/3 is that that the answer? i assumed it a/(b+c) rather than a/b + c --- which part of this do you want more explanation on?

OpenStudy (anonymous):

the original problem is written as a/b+c , I still get 7/10 that is why I asked to have it cleared up, I thought I had but I dont.

OpenStudy (asnaseer):

do you agree that it can be reduced to the following?\[\frac{a}{b+c}=\frac{5b}{b+\frac{b}{2}}\]

OpenStudy (anonymous):

thats as far as i got then i got lost.

OpenStudy (asnaseer):

ok, let me take you through it from this point on...

OpenStudy (anonymous):

please.

OpenStudy (asnaseer):

\[\frac{a}{b+c}=\frac{5b}{b+\frac{b}{2}}=\frac{5b}{\frac{2b}{2}+\frac{b}{2}}=\frac{5b}{\frac{3b}{2}}=5b\div\frac{3b}{2}=5\cancel{b}\times\frac{2}{3\cancel{b}}=\frac{10}{3}\]

OpenStudy (anonymous):

thats a little clearer I have to learn how to input fractions on this web site. thank you

OpenStudy (asnaseer):

this uses the fact that:\[x\div\frac{y}{z}=x\times\frac{z}{y}\]

OpenStudy (asnaseer):

there is a latex practising group that might help you - there are lots of tips there (mostly in the closed questions)

OpenStudy (asnaseer):

here is the link to it: http://openstudy.com/study?login#/groups/LaTeX%20Practicing!%20%3A)

OpenStudy (anonymous):

great thanks

OpenStudy (asnaseer):

yw